OpenStudy (anonymous):

solve the following equation for x 23x-17y=5

OpenStudy (anonymous):

x=17y/ 23 + 5/ 23

OpenStudy (anonymous):

Move -17y to the RHS and then divide each side by 23. Be sure to change the sign of -17y . x = (17y+5)/23 The idea is to isolate x. Prior posting a moment ago had a sign error. Sorry about that.

OpenStudy (anonymous):

im not sure if i got this right I'm really bad at math but I'm trying wouldn't it cancel out on the left side when i add 17y and when i add 17y to the right side i got 40 ??

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

23x-17y=5 23x-17y+17y=5+17y -> 23x=5+17y

OpenStudy (anonymous):

so how do we find x from there ? @robtobey

OpenStudy (anonymous):

Divide each sided of the equation by 23. 23x/23 is x (5+17y)/23 is (5+17y)/23 x = (5+17y)/23 or 5/23 + 17y/23

OpenStudy (anonymous):

im so confused @robtobey